OpenStudy (anonymous):

A right triangle is removed from a rectangle to create the shaded region shown below. Find the area of the shaded region. Be sure to include the correct unit in your answer.

OpenStudy (anonymous):



OpenStudy (anonymous):

Too little data to conclude an answer honestly :/ there has to be more :/

OpenStudy (anonymous):

I'll do a screen print just a minnute

OpenStudy (anonymous):

@Mathbreaker

OpenStudy (anonymous):

ooh... this is simple then 8-5 = 3, 6-4 = 2 so the sides of triangle are 3 and 2, now using the formula 1/2 *a * b we get 3*2/2 = 3 :)

OpenStudy (anonymous):

so the answer is 45 then?

OpenStudy (anonymous):

Yes :) Cheers :P

OpenStudy (anonymous):

ur the best

OpenStudy (anonymous):

Thanks :)